OpenStudy (anonymous):

A ball is dropped from a certain height. The function below represents the height f(n), in feet, to which the ball bounces at the nth bounce: f(n) = 9(0.7)n What does the number 0.7 represent? The ball bounces to 30% of its previous height with each bounce The height at which the ball bounces at the nth bounce is 0.3 feet The ball bounces to 70% of its previous height with each bounce The height from which the ball was dropped at the nth bounce is 0.7 feet

OpenStudy (anonymous):

@mathmate @mathlover2014

OpenStudy (mathmate):

I think you mean \(f(n) = 9(0.7)^n\) If you do not use LaTeX to write the equations, you still need to write a ^ before exponents so that others know that the following digit or letter is an exponent. You could have written f(n) = 9(0.7)^n. The physical meaning of the various parts of this exponential equation is as follows: h(x)=h0(b)^x h0=initial value, for example, height of ball before the first bounce (i.e. x=0) b=ratio of (height after each bounce) / (height before bounce) x=number of bounces.